Opened 10 years ago

Closed 6 years ago

#410 closed flight processing (fixed)

UR11_01, flight day 215/2011, Poole Burn

Reported by: knpa Owned by: knpa
Priority: alpha 4 low Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2011/flight_data/uk/UR11_01-2011_215_Poole_Burn

Data arrived from ARSF via network transfer on 04/08/2011

UR11_01 Scientific objective:
unknown

UR11_01 Priority: a4l (may increase later)

PI: Edward (Ted) J Milton

Sensors:

Camera (13/10/11)
Eagle (24/11/11)
Hawk (24/11/11)
LiDAR (24/11/11)

Change History (31)

comment:1 Changed 10 years ago by knpa

  • Summary changed from UR11_01, flight day 215/2011, Poole to UR11_01, flight day 215/2011, Poole Burn

comment:2 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Navigation processing completed.

comment:3 Changed 10 years ago by knpa

Starting LiDAR processing.

comment:4 Changed 10 years ago by knpa

las files generated.

comment:5 Changed 10 years ago by knpa

Slight horizontal discrepancy, looks like a pitch error.

comment:6 Changed 10 years ago by knpa

Have changed pitch from -0.00324419 to -0.0029 which seems to give the best result.

comment:7 Changed 10 years ago by knpa

The second line has twice the FOV of the first line (24.5 compared to 12.5).

comment:8 Changed 10 years ago by knpa

Possible flight roll for second line (~8cm error at edges). Will not correct as not enough lines to determine if this is a systematic error.

comment:9 Changed 10 years ago by knpa

Started processing HS.

comment:10 Changed 10 years ago by knpa

Started processing RCD.

Images 1-9 are non-flightline covering.

comment:11 Changed 10 years ago by knpa

RCD ready for delivery check: UR11_01-215-camera-20111004

comment:12 Changed 10 years ago by adbe

Delivery checking RCD

comment:13 Changed 10 years ago by adbe

Camera ready to deliver

comment:14 Changed 10 years ago by knpa

LiDAR is currently on hold while waiting for vectors.

comment:15 Changed 10 years ago by knpa

HS

Generated DEMs and config file.

comment:16 Changed 10 years ago by knpa

  • Description modified (diff)

Delivered RCD today (13/10/11) to Edward Milton at Southampton.

comment:17 Changed 10 years ago by knpa

APL output claims no dark frames found on either Eagle file.
DEM needs to be extended for second Hawk line.

comment:18 Changed 10 years ago by knpa

Can confirm neither of the Eagle lines have dark frames.

comment:19 Changed 10 years ago by knpa

Used dark frames from following files:

e1: VNIR244-11-15.raw
e2: VNIR165a-11-22.raw

These were scaled up to take into account slightly different integration times.

As these were taken from different projects and had this scaling applied, the intensity values are probably going to be poor compared to usual.

comment:20 Changed 10 years ago by knpa

SCTs

e h
1 -0.03 -0.03
2 -0.02 -0.06

comment:21 Changed 10 years ago by knpa

Created hyperspectral delivery. Ready for checking apart from I need to add vectors to the screenshots.

comment:22 Changed 10 years ago by knpa

Created screenshots with vectors, and Readme. Ready for deliery check at UR11_01-215-hyperspectral-20111122 (adbe to do).

comment:23 Changed 10 years ago by adbe

Hyperspectral delivery check: -

  • Mention eagle first and last bands are blank in ReadMe
  • Maybe include a misc directory in the delivery containing TN02_NTv2.gsb, so that the apl commands can be run easily. The ReadMe needs to be changed if this is done, and I need to redo an apl check.

comment:24 Changed 10 years ago by knpa

Added the band issue mention to Readme.

The TN02_NTv2.gsb file does not need to be included as this can be downloaded by the PI.

This sensor is ready for delivery, will deliver with LiDAR which is almost complete.

comment:25 Changed 10 years ago by knpa

Finishing off lidar, comparing with vectors.

comment:26 Changed 10 years ago by knpa

Lidar finished and ready for checking at: UR11_01-215-lidar-20111003

comment:27 Changed 10 years ago by adbe

LiDAR delivery check:

  • Logsheet should be named: UR11_01_2011_215_Poole_Burn.pdf. And is the original logsheet, there is one in the Hyperspectral delivery.

Otherwise, ready to deliver

comment:28 Changed 10 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

Correct above.

Delivered Hyperspectral and LiDAR today (24/11/11) to Ted Milton in Southampton.

comment:29 Changed 10 years ago by mark1

Hyperspectral XML files have been re-created and re-delivered to PI (via email). This corrects Eagle IFOV value in XML and extra comments included.

comment:30 Changed 9 years ago by knpa

Beginning archiving.

comment:31 Changed 6 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archived data available from NEODC.

Note: See TracTickets for help on using tickets.